# A Unified Framework for Analyzing Meta-algorithms in Online Convex Optimization

### v3

@@ 3 Problem Setup
${\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}}({\mathcal{U}})[a,b]$ $:=\sup_{{\mathcal{B}}\in\operatorname{Adv}}{\mathbb{E}}\left[\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\min_{{\mathbf{u}}=({\mathbf{u}}_{1},\cdots,{\mathbf{u}}_{T})\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right],$
where the expectation in the definition of the regret is over the randomness of the algorithm and the query oracles. We use the notation ${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}}({\mathcal{U}})[a,b]:={\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}}({\mathcal{U}})[a,b]$ when $\operatorname{Adv}=\{{\mathcal{B}}\}$ is a singleton.
Static adversarial regret or simply adversarial regret corresponds to $a=1$, $b=T$ and ${\mathcal{U}}={\mathcal{K}}_{\star}^{T}:=\{({\mathbf{x}},\cdots,{\mathbf{x}})\mid{\mathbf{x}}\in{\mathcal{K}}\}$. When $a=1$, $b=T$ and ${\mathcal{U}}$ contains only a single element then it is referred to as the dynamic regret (?, ?). If instead ${\mathcal{U}}={\mathcal{K}}^{T}$, then it is referred to as restricted dynamic regret (?). Adaptive regret, is defined as $\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}}({\mathcal{K}}_{\star}^{T})[a,b]$ (?). We drop $a$, $b$ and ${\mathcal{U}}$ when the statement is independent of their value or their value is clear from the context.
@@
(${\mathbb{P}}\left(\{\omega\in\Omega\mid{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}\leq h(\delta)\}\right)\geq 1-\delta.$
Note that the infimum of any family of high probability regret bounds is a high probability regret bound. Hence we use ¯$\overline{{\mathcal{R}}}_{\operatorname{Adv}}^{{\mathcal{A}}}:[0,1]\to[0,\infty)$ to denote the smallest high probability regret bound for $({\mathcal{A}},\operatorname{Adv})$. In other words, ¯$\overline{{\mathcal{R}}}_{{\mathcal{B}}}^{{\mathcal{A}}}$ is the quantile function for the random variable $\omega\mapsto{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}$.
4 Re-statement of Previous Result: Oblivious to fully adaptive adversary
The following theorem states a result that is well-known in the literature. For example, (?) only defined the notion of adaptive adversary when discussing stochastic algorithms and mentions that the results of all previous chapters, which were focused on deterministic algorithms, apply to any type of adversary. Here we explicitly mention it as a theorem for completion.
@@ 4 Re-statement of Previous Result: Oblivious to fully adaptive adversary
5 Linear to convex with fully adaptive adversary
@@
Corollary 2 (Theorem 14 in (?)).
The min-max regret bound of the problem of online linear optimization with deterministic feedback and online convex optimization with deterministic feedback are equal.
@@
6 Full information feedback to semi-bandit feedback
@@
             Return $\nabla q_{t}({\mathbf{y}}_{t,i})={\mathbf{o}}_{t}+\mu\langle{\mathbf{y}}_{t,i},{\mathbf{y}}_{t,i}-{\mathbf{x}}_{t}\rangle$ as the output of the query oracle to ${\mathcal{A}}$
@@
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}}$
@@
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}$
@@
7 Lipschitz to non-Lipschitz
It is well-known that when a convex function is restricted to a domain smaller than its original domain, then the restricted function is Lipschitz (See Observation 3 in (?) and Lemma 1). Here we use a shrinking method described in (?). We choose a point ${\mathbf{c}}\in\operatorname{relint}({\mathcal{K}})$ and a real number $r>0$ such that ${\mathbb{B}}_{r}({\mathbf{c}})\cap\operatorname{aff}({\mathcal{K}})\subseteq{\mathcal{K}}$. Then, for any shrinking parameter $0\leq\alpha<r$, we define $\hat{{\mathcal{K}}}_{\alpha}:=(1-\frac{\alpha}{r}){\mathcal{K}}+\frac{\alpha}{r}{\mathbf{c}}$. We define ${\mathbf{F}}|_{\hat{{\mathcal{K}}}_{\alpha}}$ to function class $\{f|_{\hat{{\mathcal{K}}}_{\alpha}}\mid f\in{\mathbf{F}}\}$. Given an adversary $\operatorname{Adv}$, we define $\operatorname{Adv}|_{\hat{{\mathcal{K}}}_{\alpha}}$ to be the adversary constructed by restricting the output of $\operatorname{Adv}$ to the set $\hat{{\mathcal{K}}}_{\alpha}$. Recall that the the domain ${\mathcal{K}}$ of the adversary is known to the agent ${\mathcal{A}}$. For an online algorithm ${\mathcal{A}}$, we define ${\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}$ to be the online algorithm resulting from restricting the domain of ${\mathcal{A}}$ to $\hat{{\mathcal{K}}}_{\alpha}$.
@@ 7 Lipschitz to non-Lipschitz
${\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}}({\mathcal{U}})\leq{\mathcal{R}}_{\operatorname{Adv}_{\hat{{\mathcal{K}}}_{\alpha}}}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\frac{\alpha M_{0}T}{r},\quad\overline{{\mathcal{R}}}_{\operatorname{Adv}}^{{\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}}({\mathcal{U}})\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{\hat{{\mathcal{K}}}_{\alpha}}}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\frac{\alpha M_{0}T}{r}.$
@@
8 Deterministic feedback to stochastic feedback
In this section, we show that algorithms that are designed for fully adaptive adversaries using deterministic semi-bandit feedback can obtain similar bounds using only stochastic semi-bandit feedback when facing oblivious adversaries. In particular, there is no need for any variance reduction method, such as momentum (?, ?, ?), as long as we know that the algorithm is designed for a fully adaptive adversary.
@@ 8 Deterministic feedback to stochastic feedback
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}},\quad\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}}$
See Appendix E for proof. Theorem 6.5 in (?) may be seen as a special case of the above theorem when ${\mathcal{U}}$ is a singleton and ${\mathcal{A}}$ is deterministic.
9 First order feedback to zeroth order feedback
When we do not have access to a gradient oracle, we rely on samples from a value oracle to estimate the gradient. The “smoothing trick” is a classical idea in convex optimization (See (?)) which was first used in online convex optimization in (?). This idea involves averaging through spherical sampling around a given point. Here we use a variant that was introduced in (?) which does not require extra assumptions on the convex set ${\mathcal{K}}$.
@@ 9 First order feedback to zeroth order feedback
             Let $o_{t,i}$ be the output of the query the oracle at the point ${\mathbf{y}}_{t,i}+\delta{\mathbf{v}}_{t,i}$
             Pass $\frac{k}{\delta}o_{t}{\mathbf{v}}_{t}$ as the oracle output to ${\mathcal{A}}$
@@
$\hat{f}_{\delta}({\mathbf{x}})$ $:={\mathbb{E}}_{{\mathbf{z}}\sim{\mathbb{B}}_{\delta}({\mathbf{x}})\cap\operatorname{aff}({\mathcal{K}})}[f({\mathbf{z}})]$
$={\mathbb{E}}_{{\mathbf{v}}\sim{\mathbb{B}}_{1}({\mathbf{0}})\cap{\mathcal{L}}_{0}}[f({\mathbf{x}}+\delta{\mathbf{v}})],$
@@
    ${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{{\mathcal{A}}^{\prime}}\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}(\hat{{\mathbf{F}}},\frac{k}{\delta}B_{0})}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\left(3+\frac{2D\alpha}{r\delta}\right)\delta M_{1}T.$
@@
See Appendix F for proof.
Remark 2.
Remark 3.
10 First order feedback to deterministic zeroth order feedback
When we have access to a deterministic zeroth order feedback, we may use the two point gradient estimator (?, ?) to significantly improve the results of the previous section.
Theorem 8.
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}({\mathcal{U}})\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}(\hat{{\mathbf{F}}},kM_{1})}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\left(3+\frac{2D}{r}\right)\delta M_{1}T.$
@@ 10 First order feedback to deterministic zeroth order feedback
Meta-algorithm 4 First order to zeroth order with Two Point gradient estimator - $\mathtt{FOTZO\textrm{-}2P}({\mathcal{A}})$
11 Applications
@@
Online Gradient Descent: Recall that $\mathtt{OGD}$ algorithm (?) is a deterministic algorithm designed for deterministic semi-bandit feedback that obtains a regret bound of $O\left(M_{1}T^{1/2}\right)$ over $M_{1}$-Lipschitz convex class ${\mathbf{F}}$ and $O\left(M_{1}^{2}\log(T)\right)$ over strongly convex $M_{1}$-Lipschitz convex class ${\mathbf{F}}^{\prime}$ (see (?)). Hence we may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$ and $\alpha=\delta=(T\log(T))^{-1/2}$, respectively, to see that
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\mathtt{STB}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(B_{0}T^{3/4}\right),\quad{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}}^{\prime},B_{0})}^{\mathtt{STB}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]=O\left(B_{0}^{2}T^{1/2}\log(T)^{1/2}\right).$
The first result generalizes the result of (?) to allow the feedback of the value oracle to be stochastic. The second result provides a new optimal algorithm for bandit strongly convex optimization, matching the result of (?). However, (?) obtains a regret bound of the same order for the strongly convex case while assuming access to a self-concordant barrier for the domain and also assuming that the functions are smooth. We do not require either of these assumptions here.
On the other hand, using Theorem 8 on $\mathtt{OGD}$, we see that:
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}T^{1/2}\right),\quad{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}}^{\prime})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}^{2}\log(T)\right),$
where the first result is proven in (?) and the second one is novel.
Ader and Improved Ader: In (?), Algorithm 2, i.e., “Ader”, is a full-information deterministic algorithm designed for deterministic first order feedback. Algorithm 3, i.e., “Improved Ader”, is simply the result of applying the $\mathtt{FTS}$ meta-algorithm to Ader. We may apply Theorems 5 and 7, with $\alpha=\delta=T^{-1/4}$, to see that

### v4

@@ 3. Problem Setup
${\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}}({\mathcal{U}})[a,b]$ $:=\sup_{{\mathcal{B}}\in\operatorname{Adv}}{\mathbb{E}}\left[\sum_{t=a}^{b}f_{t}({\mathbf{x}}_{t})-\min_{{\mathbf{u}}\in{\mathcal{U}}}\sum_{t=a}^{b}f_{t}({\mathbf{u}}_{t})\right]$
where ${\mathbf{u}}=({\mathbf{u}}_{1},\cdots,{\mathbf{u}}_{T})$ and the expectation in the definition of the regret is over the randomness of the algorithm and the query oracles. We use the notation ${\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}}({\mathcal{U}})[a,b]:={\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}}({\mathcal{U}})[a,b]$ when $\operatorname{Adv}=\{{\mathcal{B}}\}$ is a singleton.
Static adversarial regret or simply adversarial regret corresponds to $a=1$, $b=T$ and ${\mathcal{U}}={\mathcal{K}}_{\star}^{T}:=\{({\mathbf{x}},\cdots,{\mathbf{x}})\mid{\mathbf{x}}\in{\mathcal{K}}\}$. When $a=1$, $b=T$ and $\mathcal{U}$} contains only a single element then it is referred to as the dynamic regret zinkevich03_onlin; zhang18_adapt_onlin_learn_dynam_envir. This notion of regret allows us to provide an upper bound on regret that may depend on the comparator. In Section 11, we will discuss algorithms that provide guarantees depending on the path-length, defined with $P_{T}:={\mathbf{u}}\mapsto\sum_{t=1}^{T-1}\|{\mathbf{u}}_{t}-{\mathbf{u}}_{t+1}\|:{\mathcal{K}}^{T}\to{\mathbb{R}}$. Adaptive regret, is defined as $\max_{1\leq a\leq b\leq T}{\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}}({\mathcal{K}}_{\star}^{T})[a,b]$ hazan09_effic. A crucial point in this definition is that the comparison is with respect to a (potentially) different optimum for any interval. We drop $a$, $b$ and $\mathcal{U}$} when the statement is independent of their value or their value is clear from the context.
@@
${\mathbb{P}}\left(\{\omega\in\Omega\mid{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}\leq h(\delta)\}\right)\geq 1-\delta.$
Note that the infimum of any family of high probability regret bounds is a high probability regret bound. Hence we use $\overline{{\mathcal{R}}}_{\operatorname{Adv}}^{{\mathcal{A}}}:[0,1]\to[0,\infty)$ to denote the smallest high probability regret bound for $({\mathcal{A}},\operatorname{Adv})$. In other words, $\overline{{\mathcal{R}}}_{{\mathcal{B}}}^{{\mathcal{A}}}$ is the quantile function for the random variable $\omega\mapsto{\mathcal{R}}_{{\mathcal{B}}}^{{\mathcal{A}}_{\omega}}$.
4. Re-statement of Previous Result: Oblivious to fully adaptive adversary
The following theorem states a result that is well-known in the literature. For example, hazan2016introduction only defined the notion of adaptive adversary when discussing stochastic algorithms and mentions that the results of all previous chapters, which were focused on deterministic algorithms, apply to any type of adversary. Here we explicitly mention it as a theorem for completion.
@@ 4. Re-statement of Previous Result: Oblivious to fully adaptive adversary
5. Linear to convex with fully adaptive adversary
@@
Corollary 0 (Theorem 14 in abernethy2008optimal).
The min-max regret bounds of online linear optimization with deterministic feedback and online convex optimization with deterministic feedback are equal.
@@
6. Full information feedback to semi-bandit feedback
@@
     Return $\nabla q_{t}({\mathbf{y}}_{t,i})={\mathbf{o}}_{t}+\mu({\mathbf{y}}_{t,i}-{\mathbf{x}}_{t})$ as the output of the query oracle to $\mathcal{A}$}
@@
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}}_{\mu})}^{{\mathcal{A}}},$
@@
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{F}})}^{{\mathcal{A}}}.$
@@
7. Lipschitz to non-Lipschitz
It is well-known that when a convex function is restricted to a domain smaller than its original domain, then the restricted function is Lipschitz (See Observation 3 in flaxman2005online and Lemma 16). Here we use a shrinking method described in pedramfar23_unified_approac_maxim_contin_dr_funct. We choose a point ${\mathbf{c}}\in\operatorname{relint}({\mathcal{K}})$ and a real number $r>0$ such that ${\mathbb{B}}_{r}({\mathbf{c}})\cap\operatorname{aff}({\mathcal{K}})\subseteq{\mathcal{K}}$. Then, for any shrinking parameter $0\leq\alpha<r$, we define $\hat{{\mathcal{K}}}_{\alpha}:=(1-\frac{\alpha}{r}){\mathcal{K}}+\frac{\alpha}{r}{\mathbf{c}}$. We define ${\mathbf{F}}|_{\hat{{\mathcal{K}}}_{\alpha}}$ to function class $\{f|_{\hat{{\mathcal{K}}}_{\alpha}}\mid f\in{\mathbf{F}}\}$. Given an adversary $\operatorname{Adv}$, we define $\operatorname{Adv}|_{\hat{{\mathcal{K}}}_{\alpha}}$ to be the adversary constructed by restricting the output of $\operatorname{Adv}$ to the set $\hat{{\mathcal{K}}}_{\alpha}$. Recall that the the domain $\mathcal{K}$} of the adversary is known to the agent $\mathcal{A}$}. For an online algorithm $\mathcal{A}$}, we define ${\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}$ to be the online algorithm resulting from restricting the domain of $\mathcal{A}$} to $\hat{{\mathcal{K}}}_{\alpha}$.
@@ 7. Lipschitz to non-Lipschitz
${\mathcal{R}}_{\operatorname{Adv}}^{{\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}}({\mathcal{U}})$ $\leq{\mathcal{R}}_{\operatorname{Adv}_{\hat{{\mathcal{K}}}_{\alpha}}}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\frac{\alpha M_{0}T}{r},$
$\overline{{\mathcal{R}}}_{\operatorname{Adv}}^{{\mathcal{A}}|_{\hat{{\mathcal{K}}}_{\alpha}}}({\mathcal{U}})$ $\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{\hat{{\mathcal{K}}}_{\alpha}}}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\frac{\alpha M_{0}T}{r}.$
@@
8. Deterministic feedback to stochastic feedback
In this section, we show that algorithms that are designed for fully adaptive adversaries using deterministic semi-bandit feedback can obtain similar bounds using only stochastic semi-bandit feedback when facing oblivious adversaries. In particular, there is no need for any variance reduction method, such as momentum mokhtari20_stoch_condit_gradien_method; chen18_projec_free_onlin_optim_stoch_gradien; xie20_effic_projec_free_onlin_method, as long as we know that the algorithm is designed for a fully adaptive adversary.
@@ 8. Deterministic feedback to stochastic feedback
${\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}$ $\leq{\mathcal{R}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}},$
$\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{o}}({\mathbf{F}},B_{1})}^{{\mathcal{A}}}$ $\leq\overline{{\mathcal{R}}}_{\operatorname{Adv}_{1}^{\text{f}}({\mathbf{Q}}_{\mu}[B_{1}])}^{{\mathcal{A}}}.$
See Appendix E for proof. Theorem 6.5 in hazan2016introduction may be seen as a special case of the above theorem when $\mathcal{U}$} is a singleton and $\mathcal{A}$} is deterministic.
9. First order feedback to zeroth order feedback
@@
     Let $o_{t,i}$ be the output of the query oracle at the point ${\mathbf{y}}_{t,i}+\delta{\mathbf{v}}_{t,i}$
     Pass $\frac{k}{\delta}o_{t,i}{\mathbf{v}}_{t,i}$ as the oracle output to $\mathcal{A}$}
@@
When we do not have access to a gradient oracle, we rely on samples from a value oracle to estimate the gradient. The “smoothing trick” is a classical idea in convex optimization (See nemirovsky83_probl_compl_method_effic_optim) which was first used in online convex optimization in flaxman2005online. This idea involves averaging through spherical sampling around a given point. Here we use a variant that was introduced in pedramfar23_unified_approac_maxim_contin_dr_funct which does not require extra assumptions on the convex set $\mathcal{K}$}.
$\hat{f}_{\delta}({\mathbf{x}})$ $:={\mathbb{E}}_{{\mathbf{z}}\sim{\mathbb{B}}_{\delta}({\mathbf{x}})\cap\operatorname{aff}({\mathcal{K}})}[f({\mathbf{z}})]={\mathbb{E}}_{{\mathbf{v}}\sim{\mathbb{B}}_{1}({\mathbf{0}})\cap{\mathcal{L}}_{0}}[f({\mathbf{x}}+\delta{\mathbf{v}})],$
@@
    ${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{{\mathcal{A}}^{\prime}}({\mathcal{U}})$ $\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}(\hat{{\mathbf{F}}},\frac{k}{\delta}B_{0})}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\left(3+\frac{2D\alpha}{r\delta}\right)\delta M_{1}T,$
    where $D=\operatorname{diam}({\mathcal{K}})$.
@@
Meta-algorithm 4 First order to zeroth order with two-point gradient estimator - $\mathtt{FOTZO\textrm{-}2P}({\mathcal{A}})$
See Appendix F for proof.
Remark 0.
Remark 0.
10. First order feedback to deterministic zeroth order feedback
When we have access to a deterministic zeroth order feedback, we may use the two point gradient estimator agarwal2010optimal; shamir17_optim_algor_bandit_zero_order to significantly improve the results of the previous section.
Theorem 14.
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{{\mathcal{A}}^{\prime}}({\mathcal{U}})\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}(\hat{{\mathbf{F}}},kM_{1})}^{{\mathcal{A}}}(\hat{{\mathcal{U}}})+\left(2+\frac{2D}{r}\right)\delta M_{1}T,$
where $D=\operatorname{diam}({\mathcal{K}})$.
11. Applications
@@
Online Gradient Descent: Recall that $\mathtt{OGD}$ algorithm zinkevich03_onlin is a deterministic algorithm designed for deterministic semi-bandit feedback that obtains a regret bound of $O\left(M_{1}T^{1/2}\right)$ over $M_{1}$-Lipschitz convex class $\mathbf{F}$} and $O\left(M_{1}^{2}\log(T)\right)$ over strongly convex $M_{1}$-Lipschitz convex class $\mathbf{F}$$^{\prime}$ (see hazan2016introduction).  [^4] Here we assume $\operatorname{diam}({\mathcal{K}})\leq 1$ for simplicity. General bounds follow similarly and may depend on $\operatorname{diam}({\mathcal{K}})$ and $r$, as described in Theorems 10 and 14. Since $\mathtt{OGD}$ is deterministic, using Theorem 2, we see that
${\mathcal{R}}_{\operatorname{Adv}^{\text{f}}_{1}({\mathbf{F}})}^{\mathtt{OGD}}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}T^{1/2}\right),$
$\quad{\mathcal{R}}_{\operatorname{Adv}^{\text{f}}_{1}({\mathbf{F}}^{\prime})}^{\mathtt{OGD}}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}^{2}\log(T)\right).$
Hence we may apply Theorems 9 to see that, if we have access to stochastic gradient oracle bounded by $B_{1}\geq M_{1}$, then we have
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}({\mathbf{F}},B_{1})}^{\mathtt{OGD}}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(B_{1}T^{1/2}\right),$
$\quad{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}({\mathbf{F}}^{\prime},B_{1})}^{\mathtt{OGD}}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(B_{1}^{2}\log(T)\right).$
For convex functions, we may apply Theorem 11, with $\alpha=\delta=T^{-1/4}$. Thus, if we have access to stochastic value oracle bounded by $B_{0}\geq M_{1}$, we have
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}},B_{0})}^{\mathtt{STB}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}({\mathbf{F}},\frac{k}{\delta}B_{0})}^{\mathtt{OGD}}(\hat{{\mathcal{K}}}_{*}^{T})[1,T]+O\left(\delta M_{1}T\right)$
$=O\left(\frac{k}{\delta}B_{0}T^{1/2}+\delta M_{1}T\right)=O\left(B_{0}T^{3/4}\right),$
which generalizes the result of flaxman2005online to allow the feedback of the value oracle to be stochastic. Similarly, for strongly convex functions, we may apply Theorem 11, with $\alpha=\delta=T^{-1/3}\log(T)^{1/3}$ to see that
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}}^{\prime},B_{0})}^{\mathtt{STB}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]\leq{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{1}({\mathbf{F}}^{\prime},\frac{k}{\delta}B_{0})}^{\mathtt{OGD}}(\hat{{\mathcal{K}}}_{*}^{T})[1,T]+O\left(\delta M_{1}T\right)$
$=O\left(\frac{k^{2}}{\delta^{2}}B_{0}^{2}\log(T)+\delta M_{1}T\right)=O\left(B_{0}^{2}T^{2/3}\log(T)^{1/3}\right).$
This provides a new algorithm for bandit strongly convex optimization, matching the result of agarwal2010optimal. However, if we further assume that the functions are smooth, then the regret bound could be improved to $\widetilde{O}(\sqrt{T})$ (See hazan14_bandit_convex_optim; ito20_optim_algor_bandit_convex_optim).
On the other hand, using Theorem 14 on $\mathtt{OGD}$, we see that
${\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}T^{1/2}\right),$
$\quad{\mathcal{R}}_{\operatorname{Adv}^{\text{o}}_{0}({\mathbf{F}}^{\prime})}^{\mathtt{FOTZO\textrm{-}2P}(\mathtt{OGD})}({\mathcal{K}}_{*}^{T})[1,T]$ $=O\left(M_{1}^{2}\log(T)\right),$
where the first result is proven in shamir17_optim_algor_bandit_zero_order and the second one is novel.
Ader and Improved Ader: In zhang18_adapt_onlin_learn_dynam_envir, Algorithm 2, i.e., “Ader”, is a full-information deterministic algorithm designed for deterministic first order feedback. Algorithm 3, i.e., “Improved Ader”, is simply the result of applying the $\mathtt{FTS}$ meta-algorithm to Ader. We may apply Theorems 9 and 11, with $\alpha=\delta=T^{-1/4}$, to see that